# Hyperflux: Pruning Reveals the Importance of Weights

### v2

@@ 4.2 ImageNet-2012
5 Conclusions, Limitations and Future Work
We introduced Hyperflux, a conceptually grounded $L_{0}$ method in which we construct the notions of flux and pressure and study their relationship with weight importance. Furthermore, we postulate and validate several properties of Hyperflux that enhance its explainability. Finally, our experiments show strong performance compared to existing state-of-the-art methods.
Despite its advantages, Hyperflux has several areas which could be improved. Our method incurs at least $33\%$ of the dense network’s computational cost (see Appendix D) and demands additional hyperparameters (e.g. scheduler policy and step, $\eta_{t}$) which work well at the same values across the vision tasks we tested on, but may require adjustment on other tasks. To address some of these issues, we can treat the network sparsity as the output of a dynamical control problem and the pressure as its input, so as to tightly control the transient and steady-state sparsity $\mathcal{S}$ despite differences in the tasks. Additionally, we are interested in checking whether the empirical Neural pruning law we found generalizes to other deep learning tasks.
@@ Appendix A Analysis
A.1 Why Important Weights Generate Stronger Flux
@@
A.2 Flux Connection To The Hessian
@@
The second term, $\frac{1}{2}\,\omega_{i}^{2}\,H_{ii}^{\theta}$, contains the diagonal element, $H_{ii}^{\theta}$, of the Hessian matrix of the loss function with respect to $\theta_{i}$. This shows that our flux metric captures the linear component of the loss change, while the second term captures the quadratic component, which is generally associated with Hessian-based pruning methods like Optimal Brain Damage LeCun et al. [1989]. In Optimal Brain Damage, a weight’s saliency is estimated by $\frac{1}{2}H_{ii}\omega_{i}^{2}$, typically under the assumption that the network is at a minimum where first-order gradients are zero. On the other hand, Hyperflux prunes the weights, therefore recovering the first linear component of the Taylor expansion which becomes 0 when weights converge.
@@ B.1 Factors influencing flux $\mathcal{G}_{i}^{-}(\omega,\mathcal{T})$
Given the impact of $\eta_{\omega}$ on network convergence, we study the influence of high and low learning rates on our pruning and regrowth phases. In our experiments, we study three setups on ResNet-50 with CIFAR-10. In the first two experiments, we study how constant learning rates across the entire pruning and regrowth process affect sparsity and regrowth. We choose a high learning rate of $0.01$ and a low learning rate of $0.0001$. For our third experiment, we start with the high learning rate which is then decayed using cosine annealing to a low learning rate until the end of regrowth. For all three studies we let our scheduler guide the network towards the same sparsity rate of $1\%$. However, we observe significant differences in the regrowth stage. For the first experiment, regrowth does not occur at all, with more weights being pruned even after the pressure is set to 0, while for the low learning rate, the performance initially degrades, but is followed by a substantial regrowth stage where the number of remaining parameters increases by $60\%$. For the third experiment performance does not degrade as much as for the low learning rate and the regrowth is done in a more controlled way, experiencing an increase in remaining parameters of $35\%$. The results are illustrated in Figure 5.
@@
Figure 5: The impact of the weights’ learning rate on pruning accuracy (left) and achieved sparsity (right).
@@ Appendix C Extended experiments
C.1 Layerwise sparsity levels & Weight Histograms
@@ C.2 Implicit regrowth
Figure 10: Gradient values over time for four remaining weights in the pruned network. Blue bars show gradients when $t_{i}>0$, red when $t_{i}\leq 0$. Notice the high-magnitude red gradients (flux $\mathcal{G}_{i}^{-}(\omega,\mathcal{T})$) versus the typically smaller positive gradients (momentum $\mathcal{G}_{i}^{+}(\omega,\mathcal{T})$).
@@
Figure 11: Weight-value histograms of ResNet-50 on CIFAR-10 at two different sparsity levels. Note how the weight distribution reshapes as sparsity increases.
@@
Figure 12: Weight histograms of ResNet-50 on ImageNet during two different phases at 95.77% sparsity.
@@ Appendix D Complexity analysis
We analyze both the training-time and inference-time compute cost of Hyperflux relative to a standard dense baseline. Let $f_{d}$ denote the FLOPs required for a forward pass of the dense network, and $f_{s}$ the (reduced) FLOPs for a forward pass through the sparse weights. We approximate the backward-pass cost for sparse weights as $2f_{s}$, following common conventions, and account for the dense parameters $t$ with a full backward cost of $f_{d}$. The reason for this is that all $t$ values are updated, no matter whether their associated weight is pruned or not, thereby requiring the full cost $f_{d}$. Thus, the total training cost of Hyperflux is
@@ Appendix F Training setup and reproducibility
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main aim of the paper, as stated in the introduction and abstract, is to create a conceptually grounded, empirically strong method which is able to maintain theoretical depth with respect to pruning criteria and weight importance. We consider to have achieved this goal successfully, since we have (1) SOTA results for CIFAR-10/100 and competitive results on Imagenet (2) in depth explanations of how pruning is performed using the notions of flux and pressure, which also lead to interesting properties that we validate.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have included and discussed the limitations of our framework in the final Section 5 at the end of the paper, specifically the training overhead and potential lack of generalization when it comes to other tasks other than vision.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: While our paper does not use formal theorems or proofs, we do rely on certain mathematical derivations to infer several properties of our method. As far as we are aware, these derivations are complete and their required conditions are stated.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have included all the necessary training recipes needed for reproducing our results, including a table in the appendix with all the hyperparameters, as well as detailed algorithms for our schedulers.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Yes, we went to great lengths to ensure reproducibility and open access, including code, Readmes, a file with all the packages needed to run the code, and a config file allowing to enable/disable third parties we used like WanDb so no one will be forced to use them too. The code is uploaded as supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, we include detailed training recipes needed for practitioners.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Yes, we report appropriate information about statistical significance of the experiments, including standard deviation and mean.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes. The paper details the compute resources required (3 × NVIDIA GeForce RTX 4090 GPUs) and evaluates computational cost exclusively on ImageNet, the most comprehensive benchmark.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we conform to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Although our work falls under generic neural network optimization without direct societal impacts, the reduction in inference energy consumption is beneficial for the environment.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not have such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Standard datasets (CIFAR-10/100, ImageNet-1K) and libraries (e.g., PyTorch) are used in compliance with their respective terms. CIFAR is available for research. ImageNet is used per its terms for non-commercial research, with images subject to original copyrights.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets have been introduced.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Not applicable.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM’s were not used for any important component of this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@ 4.2 ImageNet-2012
5 Conclusions
We introduced Hyperflux, a conceptually grounded $L_{0}$ method in which we construct the notions of flux and pressure and study their relationship with weight importance. Furthermore, we postulate and validate several properties of Hyperflux, leading us to the Neural Pruning Law Hypothesis, a general property for neural pruning.
In future works, we plan on extending the experiments validating NPLH, to natural language processing, generative modelling and reinforcement learning. Furthermore, we want to test a wider range of architectures such as recurrent neural networks and transformers.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Yes. We include complexity analysis in ablation studies as well as discussions within the main body for Hyperflux.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes.
    2. (b)
      Complete proofs of all theoretical results. Yes.
    3. (c)
      Clear explanations of any assumptions. Yes,
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes. Standard datasets (CIFAR-10/100, ImageNet-1K) and libraries (e.g., PyTorch) are used in compliance with their respective terms. CIFAR is available for research. ImageNet is used per its terms for non-commercial research, with images subject to original copyrights.
    2. (b)
      The license information of the assets, if applicable. Yes.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Not Applicable.
    4. (d)
      Information about consent from data providers/curators. Not Applicable.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable.
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable.
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable.
Supplementary Material
@@ Appendix A Analysis
A.1 Hyperflux mechanics
$\sum_{h=P_{s}}^{P_{f}}(\mathcal{G}_{i}(\omega,\mathcal{T}^{h})-\frac{\gamma}{d})>0$ (7)
If and only if:
$(P_{f}-P_{s})\cdot\Bigl[\mathcal{G}_{i}^{-}\!\Bigl(\omega,\mathcal{T}^{P_{s}\to P_{f}}\Bigr)-\frac{\gamma}{d}\Bigr]>0.$ (8)
A.2 Why Important Weights Generate Stronger Flux
@@
A.3 Flux Connection To The Hessian
@@
The second term, $\frac{1}{2}\,\omega_{i}^{2}\,H_{ii}^{\theta}$, contains the diagonal element, $H_{ii}^{\theta}$, of the Hessian matrix of the loss function with respect to $\theta_{i}$. This shows that our flux metric captures the linear component of the loss change, while the second term captures the quadratic component, which is generally associated with Hessian-based pruning methods like Optimal Brain Damage LeCun et al. (1989). In Optimal Brain Damage, a weight’s saliency is estimated by $\frac{1}{2}H_{ii}\omega_{i}^{2}$, typically under the assumption that the network is at a minimum where first-order gradients are zero. On the other hand, Hyperflux prunes the weights, therefore recovering the first linear component of the Taylor expansion which becomes 0 when weights converge.
@@ B.1 Factors influencing flux $\mathcal{G}_{i}^{-}(\omega,\mathcal{T})$
Given the impact of $\eta_{\omega}$ on network convergence, we study the influence of high and low learning rates on our pruning and regrowth phases. In our experiments, we study three setups on ResNet-50 with CIFAR-10. In the first two experiments, we study how constant learning rates across the entire pruning and regrowth process affect sparsity and regrowth. We choose a high learning rate of $0.01$ and a low learning rate of $0.0001$.
@@
For our third experiment, we start with the high learning rate which is then decayed using cosine annealing to a low learning rate until the end of regrowth. For all three studies we let our scheduler guide the network towards the same sparsity rate of $1\%$. However, we observe significant differences in the regrowth stage. For the first experiment, regrowth does not occur at all, with more weights being pruned even after the pressure is set to 0, while for the low learning rate, the performance initially degrades, but is followed by a substantial regrowth stage where the number of remaining parameters increases by $60\%$. For the third experiment performance does not degrade as much as for the low learning rate and the regrowth is done in a more controlled way, experiencing an increase in remaining parameters of $35\%$. The results are illustrated in Figure 7.
@@
Figure 7: The impact of the weights learning rate on pruning accuracy (left) and achieved sparsity (right).
@@ Appendix C Extended experiments
C.1 The Neural Pruning Law Hypothesis
We present extended experiments for the Neural Prunning Law Hypothesis, validating it on ImageNet-1K dataset and ResNet-50 network, as well as VGG-19 network and Cifar-100 dataset. We see that NPLH holds for several datasets and network architectures, as well as under different optimizers and learning rates. For VGG-19, Cifar-100 we see the three same distinct regions, also observed for ResNet-50, Cifar-10. However, for ImageNet, region 1 does not appear. This is natural for an under-parameterized network, struggling to learn the dataset, since there is not much redundancy within the network and compression occurs at all densities.
All experiments for NPLH were done by setting a constant learning rate for weights over the course of training and preloading a trained network. For IMP and Taylor, we pruned 10% of the remaining weights at each 10 epochs, until accuracy of the network collapses and record the sparsity after each increase, along with the threshold value needed to prune 10% of remaining weights. We tested other types of training, such as directly increasing saliency cutoff, rather than pruning a % of weights and we observed the same results. The only difference appeared for region 3, which occurs earlier for flat saliency increase, due to its aggressivity at low densities.
For Hyperflux, we ran the setup from Property 2. We set an initial pressure in the beginning of the training and let the network converge for 300 epochs. We report the convergence sparsity produced by each pressure level and plot them, leading us to the same power-law.
(a) VGG-19, Cifar-100 on Hyperflux
(b) VGG-19, Cifar-100 on IMP
Figure 11: ResNet-50, ImageNet on IMP. Notice how region 1 does not appear, while the curvature of region 3 begins at higher sparsities, where accuracy begins degrading.
C.2 Layerwise sparsity levels & Weight Histograms
@@ C.3 Implicit regrowth
Figure 14: Gradient values over time for four remaining weights in the pruned network. Blue bars show gradients when $t_{i}>0$, red when $t_{i}\leq 0$. Notice the high‐magnitude red gradients (flux $\mathcal{G}_{i}^{-}(\omega,\mathcal{T})$) versus the typically smaller positive gradients (momentum $\mathcal{G}_{i}^{+}(\omega,\mathcal{T})$).
@@
Figure 15: Weight‐value histograms of ResNet‑50 on CIFAR‑10 at two different sparsity levels. Note how the weight distribution reshapes as sparsity increases.
@@
Figure 16: Weight histograms of ResNet‑50 on ImageNet during two different phases at 95.77% sparsity.
@@ Appendix D Complexity analysis
We analyze both the training‐time and inference‐time compute cost of Hyperflux relative to a standard dense baseline. Let $f_{d}$ denote the FLOPs required for a forward pass of the dense network, and $f_{s}$ the (reduced) FLOPs for a forward pass through the sparse weights. We approximate the backward‐pass cost for sparse weights as $2f_{s}$, following common conventions, and account for the dense parameters $t$ with a full backward cost of $f_{d}$. The reason for this is that all $t$ values are updated, no matter whether their associated weight is pruned or not, thereby requiring the full cost $f_{d}$. Thus, the total training cost of Hyperflux is